If [H+] = 1 x 10^-8, then [OH-] is_________ and the soultion is______
Answer #1

[H] = 1 × 10-8

We know [ H+ ] [oH​​​​​-​] = 10 -14

  [OH] = 10-14/10-8 = 10-6

Concentration of oh is more than the h+ hence solution is basice thank you have a nice day